OpenStudy (anonymous):

Which of the following number combinations is not allowed in an atom? A) n=2, L=1, m=0 B) n=1, L=0, m=0 C) n=4, L=3, m=-3 D) n=2, L=3, m=2

thomaster (thomaster):

There are 4 quantum numbers -the principal quantum number (n) = integer in range 1 to infinity -the angular quantum number (l) = integer in the range 0 to n-1 -the magnetic quantum number (m) = Ranges from -l to +l -and the spin quantum number (s) = either -1/2 or +1/2 So n is good in all those options. Now look at L (the angular quantum number), it's 0 to n-1. For example in A, n=2 so L = 0 to 2-1 = 0 and 1. So L is also right. now m (-l to +l) which is in this case -1, 0 or +1. So the combination in A is allowed

thomaster (thomaster):

for B, n=1, L=0 to 1-1 = 0 only. M= -0 to +0 which just stays 0 So B is also allowed

thomaster (thomaster):

Can you do the same for C?

OpenStudy (anonymous):

Hold on.

OpenStudy (anonymous):

Im not 100% sure but no?

thomaster (thomaster):

n = 4 l = 0 to 4-1 = 0 to 3 = 0, 1, 2 and 3 So n and L are right Now m?

OpenStudy (anonymous):

Where did the 0,1,2 come from? Thats where im partly confused.

thomaster (thomaster):

it's a range. when the range is 1 to 4, the numbers in this range are 1, 2 3 and 4 :)

thomaster (thomaster):

m is the range of -l to +l. So we take the biggest number of L to get the range. the biggerst number is 3 so the range is -3 to +3 the numbers are -3, -2, -1, 0, 1, 2 and 3. So C is also allowed :)

thomaster (thomaster):

Can you check if D is really the answer?

OpenStudy (anonymous):

Yes Ill let you know in like 20 minutes

thomaster (thomaster):

alright

OpenStudy (anonymous):

thank you for all your help :)

thomaster (thomaster):

No problem \(\huge\ddot\smile\)